# Quantum density estimation with density matrices: Application to quantum anomaly detection

### 22 Dec 2023 v4

@@ IV.2.2 Additional: Quantum variational circuit implementation of QAFF
Figure 3: Althought the Q-DEMDE uses classical QAFF, we present an alternative quantum implementation of QAFF, a) Hybrid QAFF strategy with quantum kernel estimation protocol [59], b) Example of a quantum circuit to prepare the QAFF with 8 Fourier features and three qubits by means of multicontrolled $R_{z}$ rotations, however, we may also prepare this state with single $R_{z}$ rotations and CNOT gates as explained in Appendix A.
Once |$\ket{\psi(\bm{x})}$ and |$\ket{\psi(\bm{y})}$ are prepared, we can compute $\hat{k}_{\{\mathbf{w}_{j}\}}(\bm{x},\bm{y})=\absolutevalue{\innerproduct{\psi(\bm{x})}{\psi(\bm{x})}{\psi(\bm{y})}}^{2}$, by the quantum kernel estimation protocol [59], see Fig. 3a. Finally, to optimize the loss function, see Eq. 11, we initialize the components of the variational parameters $\{\mathbf{w}_{j}\}$ randomly from $U[0,1]$, and use parameter-shift rule [60] to estimate the gradients.
@@ IV.3 Q-DMKDE quantum circuit: estimating expectation values of density matrices
Next, we build the unitary matrix $U^{\dagger}_{n}$, see Eq. 15, whose first quadrant is composed of the unitary matrix $V^{\dagger}$ and its fourth quadrant with an identity matrix $I$ with rank $2^{n}-d$,
@@ IV.4 Computational complexity analysis
In this section, we present the computational complexity of the proposed quantum-classical density estimation strategy Q-DEMDE using QRFF and QAFF, and we compare it with the classical density estimation algorithms KDE [14, 15] and DMKDE [27]. For this purpose, we set $T$ sample pairs to train classically the QAFF, $N$ training data points which build the training density matrix (assuming $T>N$), and $M$ test samples whose density we want to estimate, recalling that QRFF and QAFF map $D$-dimensional classical data points to a quantum feature space of dimension $d$, and $r$ is the rank of the training density matrix. Furthermore, we set $S$, the number of quantum circuit measurements to perform the density estimation with Q-DMKDE algorithm. Table 2 summarizes the parameters of the density estimation methods.
@@
Table 3: Comparison of the computational complexity of the classical and quantum density estimation methods with QRFF and QAFF in training and prediction phases, with $T$ sample pairs to train the QAFF, $N$ training data points, $M$ test samples, $D$ the dimension of the classical data, $d$ the dimension of the quantum feature map, $R$ the number of measurements to estimate the gradients of the QAFF in a quantum computer, and $S$ the number of measurements in a quantum computer to find the density estimates in the Q-DEMDE algorithm.
@@ IV.5.1 QNO-DMKDE quantum algorithm
and we conclude by performing a measurement over the first $n$ qubits, obtaining that $P(\ket{0}_{n})=(1/N)\sum_{i=0}^{N-1}\innerproduct{\psi(\mathbf{x^{*}})}{\psi(\mathbf{x}_{j})}^{2}=\bra{\psi(\mathbf{x^{*}})}\rho_{\text{train}}\ket{\psi(\mathbf{x^{*}})}$.
@@ V.1 One-dimensional quantum-classical density estimation
Figure 6: Results of one-dimensional quantum-classical density estimation with 4 Fourier features: left, Q-DEMDE with QRFF in QASM simulator, center, Q-DEMDE with QAFF learned in pennylane simulator and predictions in QASM quantum simulator, right, Q-DEMDE with QAFF learned in pennylane and Q-DMKDE predictions in real IBM-Oslo quantum computer.
We present the experiments and results of one-dimensional density estimation in both a quantum simulator and a real quantum computer.
V.1.1 Experimental setup
@@
To test the Q-DEMDE method with the quantum Fourier features, we constructed a one-dimensional probability density distribution corresponding to a mixture of two Gaussians. The training data set consisted of 1000 points sampled from the pdf, whose probability density we wanted to estimate, and the test data set consisted of 250 equidistant points in $[-7,7]$.
@@
For the prediction step, we computed the probability density estimator, see Eq. 5, which corresponds to the expected value of the training density matrix, with our proposed Q-DMKDE quantum circuit, see Fig. 4; we required 4 qubits to encode the $4\times 4$ training density matrices and we executed $12000$ shots in both the quantum simulator and the real quantum computer. Since we were working with a small number of features, we also performed a normalization of the density values by subtracting the minimum value from the density estimates; although this is not necessary for a large number of Fourier features. We evaluated the Q-DEMDE with QRFF and QAFF for density estimation, by comparing their density estimates with the real probability distribution that generated the data. For this purpose, we selected as metrics the KL divergence, MAE (mean average error) and Spearman correlation which are widely used to compare probability distributions.
@@ V.1.2 Results and discussion
Table 5: Results of 1-dimensional quantum-classical density estimation using the Q-DEMDE method with 4 QRFF and 4 QAFF in the QASM quantum simulator and the real IBM Oslo quantum computer. The best results are in bold.
We report three experiments, see Fig. 6 and Table. 5, two experiments using the noisy QASM quantum simulator of the Qiskit IBM python library for the Q-DMKDE and both QRFF and QAFF and another experiment using QAFF and computing the density estimator on the real IBM Oslo quantum computer.
In the subfigures of the left and center of Fig. 6, we show the probability density estimation with both quantum random and quantum adaptive Fourier features with the QASM simulator. In contrast to QRFF, the QAFF produces a better approximation of the pdf, especially in low-density regions. Furthermore, the density estimation metrics shown in Table 5 illustrate that QAFF is more suitable for 1-dimensional density estimation in the case of a low number of Fourier features. In the subfigure of the right of Fig. 6, we present the results of QAFF learned in a classical computer, and the density estimates predicted with the Q-DMKDE quantum algorithm in the real IBM-Oslo quantum computer. As expected, the results on the real quantum computer were much noisier than the results on the QASM simulator; however, the results and the density estimation metrics show that it is possible to perform 1-dimensional density estimation on real quantum computers with relatively good performance.
@@ V.2 Two-dimensional quantum-classical density estimation
V.2.1 Experimental setup
We estimated the probability density functions of five two-dimensional data sets sampled from five two-dimensional distributions [66], named from top to bottom: Binomial, Potential 1, Potential 2, Arc and Star Eight, see Fig. 7; the statistical characteristics of these distributions are described in the Appendix E. We performed density estimation with three models: kernel density estimation (KDE) as a reference model, and the Q-DEMDE with both QRFF and QAFF with 32 components. The training data set for the density estimation tasks corresponded to 10000 points sampled from the two-dimensional distributions, while the test data set was constructed by dividing the five two-dimensional spaces by grids of $120\times 120$ points.
We use a classical computer to build the QRFF and QAFF mappings. For each data set, we constructed the quantum random Fourier features by sampling the weights from $\mathcal{N}(\mathbf{0},\mathbf{I}_{2})$, and we used these random weights as the initial parameters to train the QAFF mapping. To illustrate the DE improvement of QAFF over QRFF, we trained the QAFF independently for each 2D data set, although we could have learned the quantum adaptive Fourier weights only once for all the 2D data sets, considering that we used the Gaussian kernel training data set for the kernel learning, see Sect. IV.2.1, which is independent of the training and test density estimation data sets. After learning the optimal weights of the QAFF, we obtained the corresponding quantum maps of the training and test data sets by adjusting the $\gamma$ parameter. For each of the five 2-dimensional data sets, we used the same $\gamma$ parameter for the KDE, and Q-DEMDE with QRFF and QAFF.
@@ V.2.2 Results and discussion
Since Q-DEMDE with QRFF and QAFF are both approximations of KDE, we aimed to replicate the 2-dimensional density estimation results of the KDE in all data sets. Despite the reduced number of quantum Fourier components, Fig. 7 and Table. 6 show that both models with adaptive and random features obtained satisfactory results for two-dimensional quantum density estimation, albeit they could have improved by increasing the number of components. Results also show that for most datasets there is an advantage of optimizing the Fourier weights with QAFF, illustrating that this adaptive mapping improves the kernel approximation. Out of the five data sets, the Arc and Star Eight were the most challenging data sets for density estimation considering that they required to approximate a kernel with a small bandwidth parameter.
@@ 32 QRFF 32 QAFF
V.2.3 Additional: Experimental results of hybrid density estimation with quantum variational QAFF
While the proposed Q-DEMDE strategy uses classical computers to optimize the quantum adaptive Fourier features, in Sect. IV.2.2, we described an alternative method to train the QAFF in quantum hardware using a quantum variational algorithm. In Appendix D, we include some additional experiments showing that it is possible to perform 2-dimensional density estimation with the Q-DEMDE method and the QAFF optimized in a Pennylane noiseless quantum simulator. Although the Q-DEMDE method with classical QAFF is a more optimal method for hybrid density estimation, the feasibility of learning the QAFF in quantum hardware may have applications to other hybrid quantum machine learning methods that, unlike the Q-DEMDE algorithm, do not require an intermediate classical computation.
@@ VI Method application: Quantum-classical anomaly detection
In this section, we state a quantum-classical anomaly detection strategy as an application of our proposed hybrid density estimation method, see Fig. 8.
Anomaly detection (AD) aims to determine which samples from a given data set are “ordinary” or “normal” (being the interpretation of “normal” defined for each particular case) and which samples depart, or are deviated, from “normal” data (commonly known as “anomalies”). Some common applications of anomaly detection methods include fraud detection [10] and medical diagnosis [68]. This problem has been widely addressed in classical computers [69, 70], while quantum anomaly detection is a more recent field, from which some proposals have shown notable speed-ups in contrast to their classical counterparts [17, 71, 72].
@@
VI.1 QAD: Experimental setup
We used the quantum-classical density estimation method Q-DEMDE to detect anomalies in a previously labeled AD data set. We used a modified version, from Ref. [73], of the Cardiotocography data set from the UCI Machine Learning Repository related to fetal heart rate. It consists of 1831 samples, where each sample contains 21 attributes, with two classes: the normal class for the inliers, and the pathologic class for the outliers. This data set was divided into three partitions: training (60%), validation (20%), and test (20%), all of which contained both normal samples and outliers in the same proportion (approximately 9.6% of all samples were marked as outliers).
As a first step, QRFF and QAFF were applied to the samples of all partitions as quantum feature mappings. We worked with 4 and 8 Fourier features corresponding to quantum states of 2 and 3 qubits respectively. The QAFF were trained using a classical computer. To train the QAFF we used the Gaussian kernel training data set, see Sect. IV.2.1, corresponding to $1\times 10^{5}$ samples from $\mathcal{N}(\mathbf{0},\frac{1}{21}\mathbf{I}_{21})$, the number of trainable parameters, which corresponds to the Fourier weights, were respectively $21\times 4=84$ and $21\times 8=168$ in the cases of 4 and 8 components. Once we learned the Fourier weights, we predicted the quantum Fourier features of the training, test and validation data sets by searching for the optimal value of $\gamma$ of the kernel (considering only powers of two ranging from $2^{-10}$ to $2^{0}$), then setting a value of $\gamma=2^{-7}$ for both QRFF and QAFF. We also used a classical computer to build the $4\times 4$ and $8\times 8$ training density matrices, see Eq. 4, and to compute their spectral decompositions. We then performed the density estimations with the 4-qubit and 6-qubit Q-DMKDE quantum algorithm in a noiseless quantum simulator of the Pennylane quantum library [67]. As it is common in anomaly detection tasks, the labels were not used to train the algorithm.
@@ VI.1 QAD: Experimental setup
We also tested the QAD algorithm on a real quantum computer. For this purpose, we first applied the QRFF and QAFF mappings to all three data partitions using a classical computer with 4 Fourier features and $\gamma=2^{-5}$. Due to satisfactory experimental results, we directly used the anomaly detection data to train the QAFF, as described in Sect. IV.2.1, instead of the Gaussian synthetic kernel training data, although for a larger number of Fourier features and more complex data sets the synthetic Gaussian data set would be the preferable choice. Once the Fourier mappings were constructed, we prepared the training density matrix from its spectral decomposition using the Q-DMKDE quantum algorithm in the real IBM-Oslo quantum computer, and estimated the probability densities of the test and validation data sets using $12000$ shots.
We report the F1 score of the “anomalous” class, the accuracy, and the AUC (area under the characteristic curve). Considering that the experiments depend on the random weights of the QRFF and the random initialization of the QAFF, we performed ten experiments with different random seeds using the Pennylane quantum simulator. In contrast, we perform only one experiment for both quantum mappings in the real IBM-Oslo quantum computer. Furthermore, we initialized the weights of the QAFF from a uniform distribution $U[0,1]$ in all experiments, considering that the results were not strongly affected by the initial Fourier weights.

### 18 Mar 2024 v5

@@ IV.2.2 Additional: Quantum variational circuit implementation of QAFF
Figure 3: Althought the Q-DEMDE uses classical QAFF, we present an alternative variational algorithm to train the QAFF in quantum hardware. (a) Hybrid QAFF strategy with quantum kernel estimation protocol [60]. (b) Example of a quantum circuit to prepare the QAFF with 8 Fourier features and three qubits by means of multicontrolled $R_{z}$ rotations, however, we may also prepare this state with single $R_{z}$ rotations and CNOT gates as explained in Appendix A.
Once |$\ket{\psi(\bm{x})}$ and |$\ket{\psi(\bm{y})}$ are prepared, we can compute $\hat{k}_{\{\mathbf{w}_{j}\}}(\bm{x},\bm{y})=\absolutevalue{\langle\psi(\bm{x})|\psi(\bm{y})\rangle}^{2}$, by the quantum kernel estimation protocol [60], see Fig. 3a. Finally, to optimize the loss function, see Eq. 11, we initialize the components of the variational parameters $\{\mathbf{w}_{j}\}$ randomly from $U[0,1]$, and use parameter-shift rule [61] to estimate the gradients.
@@ IV.3 Q-DMKDE quantum circuit: estimating expectation values of density matrices
Next, we build the unitary matrix $U^{\dagger}_{n}$ in the first $n$ qubits, see Eq. 15, whose first quadrant is composed of the unitary matrix $V^{\dagger}$ and its fourth quadrant with an identity matrix $I$ with rank $2^{n}-d$,
@@ IV.4 Computational complexity analysis
In this section, we present the computational complexity of the proposed quantum-classical density estimation strategy Q-DEMDE using QRFF and QAFF, and we compare it with the complexities of the classical density estimation algorithms KDE [14, 15] and DMKDE [27]. For this purpose, we set $T$ sample pairs to train classically the QAFF, $N$ training data points which build the training density matrix (assuming $T>N$), and $M$ test samples whose density we want to estimate, recalling that QRFF and QAFF map $D$-dimensional classical data points to a quantum feature space of dimension $d$, and $r$ is the rank of the training density matrix. Furthermore, we set $S$, the number of quantum circuit measurements to perform the density estimation with Q-DMKDE algorithm. Table 2 summarizes the parameters of the density estimation methods.
@@
Table 3: Comparison of the computational complexity of the classical and quantum density estimation methods with QRFF and QAFF in training and prediction phases, with $T$ sample pairs to train the QAFF, $N$ training data points, $M$ test samples, $D$ the dimension of the classical data, $d$ the dimension of the quantum feature map, $r$ the number of eigenvalues of the training density matrix, and $S$ the number of measurements in a quantum computer to find the density estimates with the Q-DMKDE algorithm.
@@ IV.5.1 QNO-DMKDE quantum algorithm
and we conclude by performing a measurement over the first $n$ qubits, obtaining that $P(\ket{0}_{n})=(1/N)\sum_{i=0}^{N-1}\absolutevalue{\langle\psi(\mathbf{x^{*}})|\psi(\mathbf{x}_{i})\rangle}^{2}=\bra{\psi(\mathbf{x^{*}})}\rho_{\text{train}}\ket{\psi(\mathbf{x^{*}})}$.
@@ V.1 One-dimensional quantum-classical density estimation
Figure 6: Results of one-dimensional quantum-classical density estimation with 4 Fourier features: left, Q-DEMDE with QRFF in QASM simulator, center, Q-DEMDE with QAFF learned in a classical computer and predictions in the QASM quantum simulator, right, Q-DEMDE with QAFF learned in a classical computer and Q-DMKDE predictions in real IBM-Oslo quantum computer.
We present a demonstration of one-dimensional density estimation in both a quantum simulator and a real quantum computer.
V.1.1 Setup
@@
To test the Q-DEMDE method with the quantum Fourier features, we constructed a one-dimensional probability density distribution corresponding to a mixture of two Gaussians. The training data set consisted of 1000 points sampled from the probability density function (pdf), whose probability density we wanted to estimate, and the test data set consisted of 250 equidistant points in $[-7,7]$.
@@
For the prediction step, we computed the probability density estimator, see Eq. 5, which corresponds to the expected value of the training density matrix, with our proposed Q-DMKDE quantum circuit, see Fig. 4. We used the noisy QASM quantum simulator of the Qiskit IBM python library and the real IBM-Oslo quantum computer to estimate the expected value of the $4\times 4$ training density matrix using 4 qubits and 12000 circuit measurements (shots); in Appendix E, we describe the characteristics of the IBM-Oslo quantum computer. Since we were working with a small number of features, we also performed a normalization of the density values by subtracting the minimum value from the density estimates; although this is not necessary for a large number of Fourier features. We evaluated the Q-DEMDE with QRFF and QAFF for density estimation, by comparing their density estimates with the real probability distribution that generated the data. For this purpose, we selected as metrics the Kullback-Leibler (KL) divergence, mean average error (MAE) and Spearman correlation which are widely used to compare probability distributions.
@@ V.1.2 Results and discussion
Table 5: Results of one-dimensional quantum-classical density estimation using the Q-DEMDE method with four QRFF and four QAFF in the QASM quantum simulator and the real IBM-Oslo quantum computer. The best results are shown in bold.
We report three demonstrations of density estimation, see Fig. 6 and Table. 5, two using the QASM quantum simulator for the Q-DMKDE and both QRFF and QAFF and one using QAFF and computing the density estimator on the real IBM-Oslo quantum computer.
In the subfigures of the left and center of Fig. 6, we show the probability density estimation with both quantum random and quantum adaptive Fourier features with the QASM simulator. In contrast to QRFF, the QAFF produces a better approximation of the pdf, especially in low-density regions. Furthermore, the density estimation metrics shown in Table 5 illustrate that QAFF is more suitable for 1-dimensional density estimation in the case of a low number of Fourier features. In the subfigure of the right of Fig. 6, we present the results of QAFF learned in a classical computer and the density estimates predicted with the Q-DMKDE quantum algorithm in the real IBM-Oslo quantum computer. As expected, the results on the real quantum computer were much noisier than the results on the QASM simulator; however, the results and the density estimation metrics show that it is possible to perform 1-dimensional density estimation on real quantum computers with relatively good performance.
@@ V.2 Two-dimensional quantum-classical density estimation
V.2.1 Setup
We estimated the probability density functions of five two-dimensional data sets sampled from five two-dimensional distributions [67], named from top to bottom: Binomial, Potential 1, Potential 2, Arc and Star Eight, see Fig. 7; the statistical characteristics of these distributions are described in the Appendix F. We performed density estimation with three models: kernel density estimation (KDE) as a reference model, and the Q-DEMDE with both QRFF and QAFF with 32 components. The training data set for the density estimation tasks corresponded to 10000 points sampled from the two-dimensional distributions, while the test data set was constructed by dividing the five two-dimensional spaces by grids of $120\times 120$ points.
We use a classical computer to build the QRFF and QAFF mappings. For each data set, we constructed the quantum random Fourier features by sampling the weights from $\mathcal{N}(\mathbf{0},\mathbf{I}_{2})$, and we used these random weights as the initial parameters to train the QAFF mapping. To illustrate the DE improvement of QAFF over QRFF, we trained the QAFF independently for each 2D data set, although we could have learned the quantum adaptive Fourier weights only once for all the 2D data sets, considering that we used the Gaussian kernel training data set for the kernel learning, see Sect. IV.2.1, which is independent of the training and test density estimation data sets. After learning the optimal weights of the QAFF, for each of the five 2-dimensional data sets, we set the appropriate $\gamma$ scaling parameter to predict the QRFF and QAFF; we also used the same corresponding $\gamma$ parameter for the KDE method.
@@ V.2.2 Results and discussion
Since Q-DEMDE with QRFF and QAFF are both approximations of KDE, we aimed to replicate the 2-dimensional density estimation results of the KDE in all data sets. Despite the reduced number of quantum Fourier components, Fig. 7 and Table. 6 show that both models with adaptive and random features obtained satisfactory results for two-dimensional quantum density estimation, albeit they could have been improved by increasing the number of Fourier components. Results also show that for most datasets there is an advantage of optimizing the Fourier weights with QAFF, illustrating that this adaptive mapping improves the kernel approximation. Out of the five data sets, the Arc and Star Eight were the most challenging data sets for density estimation considering that they required to approximate a kernel with a small bandwidth parameter.
@@ 32 QRFF 32 QAFF
V.2.3 Additional: Results of hybrid density estimation with quantum variational QAFF
While the proposed Q-DEMDE strategy uses classical computers to optimize the quantum adaptive Fourier features, in Sect. IV.2.2, we describe an alternative method to train the QAFF in quantum hardware using a quantum variational algorithm. In Appendix D, we include some additional quantum simulations that show that it is possible to perform 2-dimensional density estimation with the Q-DEMDE method using the QAFF optimized in a Pennylane noiseless quantum simulator. Although the Q-DEMDE method with classical QAFF has less computational complexity and it is therefore more optimal for hybrid density estimation, the feasibility of learning the QAFF mapping in quantum hardware may have applications to other hybrid quantum machine learning methods that, unlike the Q-DEMDE algorithm, do not require an intermediate classical computation.
@@ VI Method application: Quantum-classical anomaly detection
In this section, we present a quantum-classical anomaly detection strategy as an application of our proposed hybrid density estimation method, see Fig. 8.
Anomaly detection aims to determine which samples from a given data set are “ordinary” or “normal” (being the interpretation of “normal” defined for each particular case) and which samples depart, or are deviated, from “normal” data (commonly known as “anomalies”). Some common applications of anomaly detection methods include fraud detection [10] and medical diagnosis [69]. This problem has been widely addressed in classical computers [70, 71], while quantum anomaly detection is a more recent field, from which some proposals have shown notable speed-ups in contrast to their classical counterparts [17, 72, 73].
@@
VI.1 QAD: Setup
We used the quantum-classical density estimation method Q-DEMDE to detect anomalies in a previously labeled AD data set. We used a modified version, from Ref. [74], of the Cardiotocography data set from the UCI Machine Learning Repository related to fetal heart rate. It consists of 1831 samples, where each sample contains 21 attributes, with two classes: the normal class for the inliers, and the pathologic class for the outliers. This data set was divided into three partitions: training (60%), validation (20%), and test (20%), all of which contained both normal samples and outliers in the same proportion (approximately 9.6% of all samples were labeled as outliers).
As a first step, QRFF and QAFF were applied to the samples of all partitions as quantum feature mappings. We worked with 4 and 8 Fourier features corresponding to quantum states of 2 and 3 qubits respectively. The QAFF were trained using a classical computer. To train the QAFF with $\gamma=0.5$, we used the Gaussian kernel training data set, see Sect. IV.2.1, corresponding to $1\times 10^{5}$ samples from $\mathcal{N}(\mathbf{0},\frac{1}{21}\mathbf{I}_{21})$, the number of trainable parameters, which corresponds to the Fourier weights, were respectively $21\times 4=84$ and $21\times 8=168$ in the cases of 4 and 8 components. Once we learned the Fourier weights, we predicted the quantum Fourier features of the training, test and validation data sets by searching for the optimal value of $\gamma$ of the kernel (considering only powers of two ranging from $2^{-10}$ to $2^{0}$), then setting a value of $\gamma=2^{-7}$ for both QRFF and QAFF. We also used a classical computer to build the $4\times 4$ and $8\times 8$ training density matrices, see Eq. 4, and to compute their spectral decompositions. We then performed the density estimations with the 4-qubit and 6-qubit Q-DMKDE quantum algorithm in a noiseless quantum simulator of the Pennylane quantum library [68]. As it is common in anomaly detection tasks, the labels were not used to train the algorithm.
@@ VI.1 QAD: Setup
We also performed a demonstration of the QAD algorithm on the real IBM-Oslo quantum computer, refer to Appendix E for its characteristics. To achieve this, we first applied the QRFF and QAFF mappings to all three data partitions using a classical computer with 4 Fourier features and $\gamma=2^{-5}$. Due to satisfactory results, we directly used the anomaly detection data to train the QAFF, as described in Sect. IV.2.1, instead of the Gaussian synthetic kernel training data, although for a larger number of Fourier features and more complex data sets the synthetic Gaussian data set would be the preferable choice. Once the Fourier mappings were constructed, we used 4 qubits to prepare the training density matrix from its spectral decomposition using the Q-DMKDE quantum algorithm in the IBM-Oslo quantum processor, and estimated the probability densities of the test and validation data sets using $12000$ shots.
We report the F1 score of the “anomalous” class, the accuracy, and the AUC (area under the characteristic curve). Since the random weights of the QRFF and the random initialization of the QAFF produce slightly different DE results in the case of a small number of Fourier features, we performed ten trials with different random seeds using the Pennylane quantum simulator. In contrast, in the real IBM-Oslo quantum computer, we performed the quantum demonstration only once for both quantum mappings. Furthermore, we initialized the weights of the QAFF from a uniform distribution $U[0,1]$ in all cases, considering that the results were not strongly affected by the initial distribution of the Fourier weights.